# From Friendship Networks to Classroom Dynamics: Leveraging Neural Networks, Instrumental Variable and Genetic Algorithms for Optimal Educational Outcomes

### v4

@@ 4.2 Instrumental variable peer effect measurement
The existing peer effect literature largely adopts the following linear-in-means specification:
$\underbrace{y_{ics}}_{\text{Student i's 8th-grade cognitive test score}}=\beta\tilde{W}_{ics}+X_{ics}\gamma+\underbrace{\theta_{s}}_{\text{Fixed Effect (FE)}}+\underbrace{\mu_{cs}}_{\text{Random Effect (RE)}}+\epsilon_{ics}$ (2)
where $y_{ics}$ is cognitive ability test score of student $i$ from class $c$ school $s$, $\tilde{W}_{ics}=\frac{\sum_{j=1,j\neq i}^{N_{cs}}z_{jcs}}{N_{cs}-1}$ is student $i$’s middle school classmates’ average 6th-grade class quantile, $X_{ics}$ is a set of control variables for student $i$ such as the student’s 6th-grade class quantile, gender, race, and parent’s education level. $\theta_{s}$ and $\mu_{cs}$ are the school fixed effect and classroom random effect, respectively. We can rewrite linear-in-means model as the $y_{cs}=\beta W_{cs}z_{cs}+X_{cs}\gamma+\theta_{s}+\mu_{cs}+\epsilon_{ics}$ where
@@
The linear-in-means model implicitly assumes that all students exert the same amount of peer influence on all their classmates as evidenced by the $W_{cs}$ matrix which has column sums equal to 1 for each column. This rules out the possibility of heterogeneous influence among all classmates and hence, no matter how we formulate classroom assignments, all students have the same total influence on all their classmates (sum up to 1) and there is no strategic scope to optimize classroom assignments. Eq(3) deviates from this setup.
Taking the $\Omega$ modeled by PeerNN as the true adjacency-probability matrix, we replace the linear-in-means specification with friendship-weighted specification, Eq(3), generating a more meaningful interpretation of $\beta$ for policymaker (See Appendix E for detail of linear-in-means model and its comparison to Eq(3)). However, we face an endogeneity issue with the $\Omega$ matrix which we illustrate with example C.4. Student $i$’s friendship selection, ${\Omega}_{ics}$, is correlated with some unobserved characteristics of student $i$ in $\epsilon_{ics}$.
$y_{ics}=\beta\tilde{\Omega}_{ics}+X_{ics}\gamma+\theta_{s}+\mu_{cs}+\epsilon_{ics}$ (3)
@@
We resort to the classical instrumental variable (IV) approach to address the endogeneity. The instrument that we use is the average 6th-grade class quantile of middle school classmates, $\tilde{W}$ in Eq(2). A valid instrument needs to fulfill two properties: exclusion constraint and relevance constraint.
@@
Exclusion constraint requires the instrument (average 6th-grade class rank for all classmates) to be uncorrelated with the unobserved confounder. This requirement is trivially satisfied since all classroom assignments are random conditional on school choice which is a covariate included in Eq(3). Hence, any confounder that is individual $i$-specific and predetermined before middle school should not correlate with the average characteristics of the classmates that are randomly assigned to him. Relevance constraint requires the instrument and the endogenous variable to be correlated. Both average 6th-grade class rank ($W_{cs}z_{cs}$) and friends’ weighted average 6th-grade class rank ($\Omega_{cs}z_{cs}$) are dependent on all classmates’ 6th-grade class rank ($z_{cs}$). Therefore, the relevance constraint is satisfied. We exemplify our explanations why both properties are satisfied with Example C.5.
@@
We estimate Eq(3) using a two-stage method that resembles the classical two-stage least-square procedure:
  1. 1.
    Run OLS for reduced form equation (4) and get fitted values of $\tilde{\Omega}$.
@@
    $\tilde{\Omega}_{ics}=\pi_{1}\underbrace{\frac{\sum_{j=1,j\neq i}^{N_{cs}}z_{jcs}}{N_{cs}-1}}_{\text{Classmates' average 6th-grade ranking}}+X_{ics}\delta+\phi_{s}+\eta_{ics}$ (4)
  2. 2.
    Substitute $\tilde{\Omega}$ in Eq(3) with the fitted value and run MLE assuming normally distributed $\mu$ and $\epsilon$.
@@
We also implement other standard methods for IV and random effect (Balestra and Varadharajan-Krishnakumar, 1987; Baltagi, 1981; Han, 2016) to check the robustness of the regression results.
@@
The two $\beta$s in Eq(2) and Eq(3) have different interpretations. For linear-in-means model, $\beta$ in Eq(2) is the marginal effect of middle school classmates’ average 6th-grade class quantile on the outcome; for friendship-weighted specification, $\beta$ in Eq(3) is the marginal effect of middle school friends’ weighted average 6th-grade class quantile on the outcome. [^1] Both marginal effect parameters (the two $\beta$s) reflect causality: the former comes from a randomized experiment whereas the latter uses IV to establish causality. The former has very little policy implication to classroom assignment since the linear-in-means model essentially restricts how much total peer effect each student can exert on other classmates (See detail in Appendix E). Our friendship-weighted specification relaxes this restriction: a popular student is allowed to exert more influence on all his classmates than another less popular student. This allows us to devise an average-maximizing class assignment if $\beta$ in Eq(3) is estimated to be statistically significant.
@@ 4.3.1 Genetic algorithm (GA)
In this section, we illustrate some specific features of the GA design that we use to optimize class assignment policy. We elaborate on the GA design that we use with Algorithm 1. Some of the referenced functions in Algorithm 1 are detailed in Appendix F. We summarize GA’s essence for those interested in the general concept of GA. Step 0: After a random initialization, GA evaluates the fitness of the initialized classroom assignment based on a fitness function (either Eq(5) or Eq(6)). Step 1: GA randomly swaps a pair of students to generate a new classroom assignment and evaluates its fitness. Step 2: GA repeats Step 1 with a new pair of students swapped each time. The number of repetitions is user-specified. Step 3: GA selects the best classroom assignment among those evaluated in Steps 0-2 and updates the optimal classroom assignment as that. Step 4: Repeat Steps 0-3 with the updated optimal classroom assignment as the initial assignment until some stopping criterion is satisfied. Note that this is a schematic summary of GA and many modifications are made when we apply GA. Readers should refer to Algorithm 1 for the exact implementation of our GA (and AFGA) design.
Algorithm 1 GA used to optimize class assignment policy
Input: $C$triangleright Set $C$ consists of all students the principal needs to divide into two classrooms
$N_{b}\leftarrow\sum_{s\in C}\mathbbm{1}_{\{s\text{ is boy}\}}$
$N_{g}\leftarrow\sum_{s\in C}\mathbbm{1}_{\{s\text{ is girl}\}}$
$N\leftarrow N_{b}+N_{g}$triangleright Alternatively, $N=\lvert C\rvert$
$j=\operatorname*{\arg\!\min}_{j}\{N_{j}\}_{j\in\{b,g\}}$
do
    $C_{1}\leftarrow C[v]$
    $C_{1}\leftarrow C\backslash C_{1}$triangleright This initialization ensures that $-1\leq\lvert C_{1}\rvert-\lvert C_{2}\rvert\leq 1$
while check$(C_{1},C_{2},N_{b},N_{g},j)=0$triangleright See Algorithm 2 in Appendix F
L = 150
M = 100
$PH={\rm matrix}(0,L,N)$triangleright Record all the chosen class assignment policies
$FSH={\rm rep}(0,L)$triangleright Record all fitness scores of the chosen policies
for $l$ in $1:L$ do
    $b\leftarrow B\sim{\rm Bernoulli}(0.05)$
    if b = 1 then $\triangleright$ Mutation starts
         do
             $CP\leftarrow(C_{1}[t_{1}],C_{2}[t_{2}])$
             $C_{1}^{\prime},C_{2}^{\prime}\leftarrow{\rm swap}(C_{1},C_{2},CP)$triangleright See Algorithm 3 in Appendix F
         while check$(C_{1}^{\prime},C_{2}^{\prime},N_{b},N_{g},j)=0$triangleright Mutation ends
    $\triangleright$ If not mutation, then crossover
         $FS={\rm rep}(0,M)$
         $CS\leftarrow(C_{1}[v_{1}],C_{2}[v_{2}])$triangleright $CS$ denotes M candidate swaps
         for $m\in 1:M$ do $\triangleright$ Crossover loop
             $(C_{1}^{\prime},C_{2}^{\prime})={\rm swap}(C_{1},C_{2},CS_{m})$triangleright $CS_{m}$ is the $m$th candidate swap
             if check$(C_{1}^{\prime},C_{2}^{\prime},N_{b},N_{g},j)=1$ then
                 $FS[m]\leftarrow{\rm fit}(C_{1}^{\prime},C_{2}^{\prime})$triangleright ‘fit’ is either Eq(5) or Eq(6)
             else
                 $FS[m]\leftarrow-\infty$
             end if
         end for
         $bp\leftarrow\operatorname*{\arg\!\max}_{m}FS[m]$
         if $FS[bp]>FSH[l-1]$ then $\triangleright$ Better than last round chosen policy
             $CP\leftarrow CS_{bp}$
         else
             $CP\leftarrow({\rm NULL},{\rm NULL})$
         end $\triangleright$ Crossover ends
    end if
    $(C_{1},C_{2})\leftarrow{\rm swap}(C_{1},C_{2},CP)$
    $PH[l,]=(C_{1},C_{2})$
    $FSH[l]={\rm fit}(C_{1},C_{2})$
end for
$bp\leftarrow\operatorname*{\arg\!\max}_{l}FSH[l]$
$C_{1},C_{2}\leftarrow PH[bp,]$
@@
We use a selection-mutation-crossover-termination four-step design. One special feature of our GA design is the incorporation of gender-ratio constraint: a swap that violates this gender ratio constraint will not be permitted. Given that all schools in our sample have close to a 50% gender ratio, multitudes of candidate policies that satisfy the gender ratio requirement are guaranteed to exist. Additionally, we constrain that the classrooms’ sizes differ by at most 1, ensuring the classrooms are of comparable sizes. See Appendix F for a detailed description of the institutional constraint and the parameter values that we set for GA.
4.3.2 Fitness function
@@
The most natural fitness function for a principal is the mean predicted peer effect for all students, $\frac{1}{\sum_{c=\{1,2\}}N_{c}}\sum_{c=\{1,2\}}\beta\mathbf{1_{N_{c}}^{\top}}\tilde{\Omega}_{c}$, where $\tilde{\Omega}_{c}=\Omega_{c}z_{c}$ is a vector of each student’s friends’ 6th-grade average class quantile. This fitness function specification potentially causes algorithmic fairness problems. GA may maximize the average at the expense of certain small groups of students’ educational outcomes. We actively combat this problem by penalizing both the standard errors of peer effect within classrooms ${\rm SE}_{\{\beta\tilde{\Omega}_{1}\}}$ and ${\rm SE}_{\{\beta\tilde{\Omega}_{2}\}}$ as well as standard error of peer effect across classrooms ${\rm SE}_{\{\beta\tilde{\Omega}\}}$, where $\tilde{\Omega}=(\tilde{\Omega}_{1},\tilde{\Omega}_{2})$. The two fitness functions are given as follows:
${\rm fit}_{{\rm GA}}(C_{1},C_{2})=$ $\frac{1}{\sum_{c=\{1,2\}}N_{c}}\sum_{c=\{1,2\}}\beta\mathbf{1_{N_{c}}^{\top}}\tilde{\Omega}_{c}$ (5)
${\rm fit}_{{\rm AFGA}}(C_{1},C_{2})=$ $\frac{1}{\sum_{c=\{1,2\}}N_{c}}\sum_{c=\{1,2\}}\beta\mathbf{1_{N_{c}}^{\top}}\tilde{\Omega}_{c}-\sum_{c=\{1,2\}}\phi{\rm SE}_{\{\beta\tilde{\Omega}_{c}\}}-\rho{\rm SE}_{\{\beta\tilde{\Omega}\}}.$ (6)
@@
where $C_{1}$ and $C_{2}$ denote the set of students assigned to the two classrooms, respectively. Note that $(C_{1},C_{2})$ fully pins down the class assignment policy, $\pi$. The penalties are weighted by $\phi$ and $\rho$.
5 Results
@@ 5.1 Friendship formation pattern
Figure 2: Heat map for classroom 1’s adjacency-probability matrix $\Omega$. The $(i,j)$-th entry, $\Omega_{ij}$, denotes the probability of student $i$ considering student $j$ as best friend. Students are sorted by gender. In Appendix G, we show that the friendship formation patterns reflected in this figure is also true in other classrooms.
@@
PeerNN can capture the most salient features in the classroom network. First, it predicts extremely strong gender homophily. We reorder all students by gender in the first classroom in our dataset and plot a heat map of $\Omega$, the adjacency-probability matrix (See Figure 2). Two dark diagonal blocks indicate that students are more likely to make friends of the same gender as themselves. Second, PeerNN finds some popular individual students, a common observation in social networks. Many social networks exhibit centrality property where certain nodes are particularly well-connected with other nodes. Third, PeerNN is, to some extent, able to capture clustering within gender groups. We observe from Figure 2 that each student has heterogeneous popularity across classmates of the same gender group, a necessary condition implied by clustering.
@@
We compare PeerNN with the conventional linear-in-means model. CEPS asks students to provide ARD responses about their friends in terms of ten distinct traits. We use $\Omega$ to predict friendship formation for all students and compare our prediction with the ARD responses. How we compute prediction error is detailed in Algorithm 4 in Appendix G. In Figure 3, we present violin plots of prediction error. For half of the ten traits including gender, local/nonlocal, smoke or drink, our model predicts better for all 1000 rounds. PeerNN also predicts better than the linear-in-means model for the rest of the traits except for ‘hardworking’. It could be due to the fact students’ own evaluation of whether they are hardworking systematically differs from their friends’ evaluation of them. Nevertheless, Overall, our model beats the linear-in-means model in terms of out-of-sample prediction by a huge margin.
@@
Figure 3: PeerNN prediction performance full version. Blue and yellow color distributions represent the empirical densities of PeerNN and linear-in-means models’ prediction errors, respectively. PeerNN outperforms the linear-in-means model in nine out of ten traits. The only trait that PeerNN performs worse than the linear-in-means model. It is likely due to how students evaluate whether they are hardworking or not may systematically differ from how they evaluate whether their friends are hardworking or not. Also, note that the scale for the ‘hardworking’ trait is much smaller than some of the other traits such as gender and local.
We present all ten violin plots in Figure 3. CEPS asks students to provide information about their friends in terms of ten distinct traits. We randomly draw without replacement $B_{i}$ friends for a student based on the multinomial distribution $\Omega_{i}$, the $ith$ row of $\Omega$, predict each of the traits, and record the prediction error for each trait. We repeat this step 1000 times for each student. For half of the ten traits including gender, local/nonlocal, smoke or drink, our model predicts better for all 1000 rounds. For the rest of the traits, our model prediction is better than the linear-in-means model most of the time except for the ‘hardworking’ trait. Overall, our model beats the linear-in-means model in terms of out-of-sample prediction by a huge margin.
5.2 Positive peer effect from ‘good students’

### v5

@@ 3.2 Instrumental variable peer effect measurement
3.2.1 Incorporating network information into linear-in-means
The existing peer effect literature largely adopts a linear-in-means specification:
$\underbrace{y_{ics}}_{\text{Student i's 8th-grade cognitive test score}}=\beta\tilde{W}_{ics}+X_{ics}\gamma+\underbrace{\theta_{s}}_{\text{Fixed Effect (FE)}}+\underbrace{\mu_{cs}}_{\text{Random Effect (RE)}}+\epsilon_{ics}$ (3.2)
where $y_{ics}$ is cognitive ability test score of student $i$ from class $c$ school $s$, $\tilde{W}_{ics}=\frac{\sum_{j=1,j\neq i}^{N_{cs}}z_{jcs}}{N_{cs}-1}$ is student $i$’s middle school classmates’ average 6th-grade class quantile, $X_{ics}$ is a set of control variables for student $i$ such as the student’s 6th-grade class quantile, gender, race, and parents’ education level. $\theta_{s}$ and $\mu_{cs}$ are the school fixed effect and classroom random effect, respectively. We can rewrite linear-in-means model as the $y_{cs}=\beta W_{cs}z_{cs}+X_{cs}\gamma+\theta_{s}+\mu_{cs}+\epsilon_{cs}$ where
@@ 3.2.1 Incorporating network information into linear-in-means
Under our framework, the linear-in-means model implicitly assumes that all students exert the same amount of peer influence on all their classmates as evidenced by the $W_{cs}$ matrix which has column sums equal to 1 for each column. This rules out the possibility of heterogeneous influence among all classmates and hence, no matter how we formulate classroom assignments, all students have the same total influence on all their classmates (sum up to 1) and there is no strategic scope to optimize classroom assignments. Our specification (3.3) deviates from this setup.
Taking the $\Omega$ modeled by PeerNN as the true friendship intensity matrix, we replace the linear-in-means specification with a friendship-weighted specification (3.3), generating a more meaningful interpretation of $\beta$ for policymaker. The $\beta$ in (3.2) measures peer effect from an average classmate, whereas $\beta$ in (3.3) measures the peer effect from an average friend. (See Appendix D for detail of linear-in-means model and its comparison to (3.3)).
$y_{ics}=\beta\tilde{\Omega}_{ics}+X_{ics}\gamma+\theta_{s}+\mu_{cs}+\epsilon_{ics}$ (3.3)
@@ 3.2.2 Addressing endogeneity with IV
We face an endogeneity issue with the $\Omega$ matrix, which we illustrate with example B.3 and depict with the orange components in Figure 4. Student $i$’s friendship selection, ${\Omega}_{ics}$, is correlated with some unobserved characteristics of student $i$ in $\epsilon_{ics}$. $\epsilon_{ics}$ can also directly impact student $i$’s cognitive test score. We resort to the classical instrumental variable (IV) approach to address the endogeneity. The instrument that we use is the average 6th-grade class quantile of middle school classmates, $\tilde{W}$ in (3.2). A valid instrument needs to fulfill two properties: exclusion constraint and relevance constraint.
Figure 4: Directed acyclic graph for the illustration of causality. The orange components depict the endogeneity problem, which we address in Section 3.2.2. The green components are the potentially correlated mechanisms which are addressed in Section 3.2.3.
The relevance constraint requires the instrument and the endogenous variable to be correlated. Both average 6th-grade class rank ($W_{cs}z_{cs}$) and friends’ weighted average 6th-grade class rank ($\Omega_{cs}z_{cs}$) are dependent on all classmates’ 6th-grade class rank ($z_{cs}$). Therefore, the relevance constraint is satisfied. In Figure 4, the relevance constraint is represented by the arrow pointing from $\tilde{W}$ to $\tilde{\Omega}$.
@@
The exclusion constraint requires the instrument (average 6th-grade class rank for all classmates) to be uncorrelated with the unobserved confounder. Barring the correlated mechanisms shown as the green component in Figure 4, this requirement is satisfied in theory since all classroom assignments are random conditional on school choice, which is a covariate included in (3.3). Hence, any confounder that is individual $i$-specific and predetermined before middle school should not correlate with the average characteristics of the classmates that are randomly assigned to him. We will address the correlated mechanisms (green components) separately in Section 3.2.3.
@@
We estimate (3.3) using a two-stage method that resembles the classical two-stage least-squares procedure. First, run OLS for the reduced form equation (3.4) and get fitted values of $\tilde{\Omega}$.
@@
$\tilde{\Omega}_{ics}=\pi_{1}\underbrace{\frac{\sum_{j=1,j\neq i}^{N_{cs}}z_{jcs}}{N_{cs}-1}}_{\text{Classmates' average 6th-grade ranking}}+X_{ics}\delta+\phi_{s}+\eta_{ics}$ (3.4)
@@
Second, substitute $\tilde{\Omega}$ in (3.3) with the fitted value and run MLE assuming normal $\mu$ and $\epsilon$. We also implement other standard methods for IV and random effect (Balestra and Varadharajan-Krishnakumar, 1987; Baltagi, 1981; Han, 2016) to check the robustness of the regression results.
3.2.3 Addressing correlated mechanisms
While our paper focuses on the peer effect that operates through friendships $\Omega$ (and hence, $\tilde{\Omega}$), there are other correlated mechanisms that do not operate through friendship, e.g., competitiveness of the classroom environment, quality of classroom discussion, and teachers’ adjustments for the level of instruction. In these cases, peer effect can exist between two students even when they are not friends, i.e., $z$ has a direct impact on $y$ without going through $\tilde{\Omega}$. Consequently, all classmates’ 6th-grade class rank $z$ becomes a confounder, as shown by Figure 4. To address this concern, we augment the base specifications with classroom-level covariates to close the backdoor path depicted in green. These control variables include the information about the 6th-grade class ranks of the entire classroom including the student themselves (25th, 75th quantile of $z_{c}$ and whether $z_{ics}$ is above the median of $z_{c}$). The rationale behind this augmentation is that the correlated mechanisms are likely functions of these additional control variables. For example, how competitive a classroom is likely a function of top students’ performance (25th quantile $z_{c}$), whereas teachers’ instruction pace is more likely a function of the bottom students’ performance (75th quantile of $z_{c}$).
Including other functions of $z_{c}$, e.g., quantiles of $z_{c}$, into the specifications potentially leads to collinearity problems, which may make our estimate not significant when there is a true peer effect. Hence, we conduct further robustness checks by replacing the quantile of $z_{c}$ with the quantile of 6th-grade persistence measures. The persistence measures close the backdoor path similarly to the quantiles of $z_{c}$: classroom competitiveness depends on how persistent students are, and teacher adjust their teaching pace based on how persistent students are.
3.2.4 Policy significance
@@
The two $\beta$s in (3.2) and (3.3) have different interpretations. For linear-in-means model, $\beta$ in (3.2) is the marginal effect of middle school classmates’ average 6th-grade class quantile on the outcome; for friendship-weighted specification, $\beta$ in (3.3) is the marginal effect of middle school friends’ weighted average 6th-grade class quantile on the outcome. [^1] Both marginal effect parameters (the two $\beta$s) reflect causality: the former comes from a randomized experiment whereas the latter uses IV to establish causality. The former has very little policy implication to classroom assignment since the linear-in-means model essentially restricts how much total peer effect each student can exert on other classmates (See detail in Appendix D). Our friendship-weighted specification relaxes this restriction: a popular student is allowed to exert more influence on all his classmates than another less popular student. This allows us to devise an average-maximizing class assignment if $\beta$ in (3.3) is estimated to be statistically significant.
@@ 3.3.1 Genetic algorithm (GA)
In this section, we illustrate some specific features of the GA design that we use to optimize classroom assignment policy. We elaborate on the GA design that we use with Algorithm 3, which is presented in Appendix E. We summarize GA’s essence for those interested in the general concept of GA as follows:
  • Step 0: Initialize a random classroom assignment that satisfies the gender ratio constraint.
  • Step 1: Draw a Bernoulli variable, if it is 1, then randomly pick a pair of students, one from each classroom, and swap them, evaluate the new classroom assignment’s fitness like in Step 0, repeat step 1. Else if the Bernoulli variable is 0, proceed to step 2.
  • Step 2: GA randomly swaps a pair of students to generate a new classroom assignment and evaluates its fitness.
  • Step 3: GA repeats Step 2 with a new pair of students swapped each time. The number of repetitions is user-specified.
  • Step 4: GA selects the best classroom assignment among those evaluated in Steps 1-3 and updates the optimal classroom assignment as that.
  • Step 5: Repeat Steps 1-4 with the updated optimal classroom assignment as the initial assignment until some stopping criterion is satisfied.
Note that this is a schematic summary of GA and many modifications are made when we apply GA. Readers should refer to Algorithm 3 for the exact implementation of our GA (and AFGA) design.
@@
We use a selection-mutation-crossover-termination four-step design. One special feature of our GA design is the incorporation of gender-ratio constraint: a swap that violates this gender ratio constraint will not be permitted. Given that all schools in our sample have close to a 50% gender ratio, multitudes of candidate policies that satisfy the gender ratio requirement are guaranteed to exist. Additionally, we constrain that the classrooms’ sizes differ by at most 1, ensuring the classrooms are of comparable sizes. See Appendix E for a detailed description of the institutional constraint and the parameter values that we set for GA.
3.3.2 Fitness function and fairness consideration
@@
Our schematic summary of GA picks the average peer effect as the fitness function used to evaluate a policy. Mathematically, the mean predicted peer effect for all students is written as $\frac{1}{\sum_{c=\{1,2\}}N_{c}}\sum_{c=\{1,2\}}\beta\mathbf{1_{N_{c}}^{\top}}\tilde{\Omega}_{c}$, where $\tilde{\Omega}_{c}=\Omega_{c}z_{c}$ is a vector of each student’s friends’ 6th-grade average class quantile. This fitness function specification potentially causes algorithmic fairness problems. GA may maximize the average at the expense of certain small groups of students’ educational outcomes. We actively combat this problem by penalizing both the standard errors of peer effect within classrooms ${\rm SE}_{\{\beta\tilde{\Omega}_{1}\}}$ and ${\rm SE}_{\{\beta\tilde{\Omega}_{2}\}}$ as well as standard error of peer effect across classrooms ${\rm SE}_{\{\beta\tilde{\Omega}\}}$, where $\tilde{\Omega}=(\tilde{\Omega}_{1},\tilde{\Omega}_{2})$. The two fitness functions are given as follows:
${\rm fit}_{{\rm GA}}(C_{1},C_{2})=$ $\frac{1}{\sum_{c=\{1,2\}}N_{c}}\sum_{c=\{1,2\}}\beta\mathbf{1_{N_{c}}^{\top}}\tilde{\Omega}_{c}$ (3.5)
${\rm fit}_{{\rm AFGA}}(C_{1},C_{2})=$ $\frac{1}{\sum_{c=\{1,2\}}N_{c}}\sum_{c=\{1,2\}}\beta\mathbf{1_{N_{c}}^{\top}}\tilde{\Omega}_{c}-\sum_{c=\{1,2\}}\phi{\rm SE}_{\{\beta\tilde{\Omega}_{c}\}}-\rho{\rm SE}_{\{\beta\tilde{\Omega}\}}.$ (3.6)
@@
where $C_{1}$ and $C_{2}$ denote the set of students assigned to the two classrooms, respectively. Note that $(C_{1},C_{2})$ fully pins down the class assignment policy, $\pi$. The penalties are weighted by $\phi$ and $\rho$. Throughout the paper, we set $\phi=\rho$.
4 CEPS data fitting results
We implement our three-step empirical framework with CEPS data. This section outlines the key findings of our application.
@@ 4.1 Friendship formation pattern
PeerNN can capture the most salient features in the classroom network. First, it predicts extremely strong gender homophily. We reorder all students by gender in the first classroom in our dataset and plot a heat map of $\Omega$, the adjacency-probability matrix (See Figure 5). Two dark diagonal blocks indicate that students are more likely to make friends of the same gender as themselves. Second, PeerNN finds some popular individual students as indicated by the dark columns. These students are the central nodes of the classroom networks. Many social networks exhibit a centrality property where certain nodes are particularly well-connected with other nodes. In Step 3, GA will attempt to make good students the central nodes.
Figure 5: Heat maps for two classrooms to demonstrate that PeerNN prediction aligns with known property of social network: (1) gender homophily (2) presence of central nodes
@@
We compare PeerNN with the conventional linear-in-means model. CEPS asks students to provide ARD responses about their friends in terms of ten distinct traits. We use $\Omega$ to predict friendship formation for all students and compare our prediction with the ARD responses. How we compute prediction error is detailed in Algorithm 4 in Appendix F. In Figure 6, we present violin plots of prediction error. For half of the ten traits including gender, local/nonlocal, smoke or drink, our model predicts better for all 1000 rounds. PeerNN also predicts better than the linear-in-means model for the rest of the traits except for ‘hardworking’. It could be due to the fact students’ own evaluation of whether they are hardworking systematically differs from their friends’ evaluation of them. Nevertheless, Overall, our model beats the linear-in-means model in terms of out-of-sample prediction by a huge margin.
@@
Figure 6: PeerNN prediction performance full version. Blue and orange color distributions represent the empirical densities of PeerNN and linear-in-means models’ prediction errors, respectively. PeerNN outperforms the linear-in-means model in nine out of ten traits. The only trait that PeerNN performs worse than the linear-in-means model. It is likely due to how students evaluate whether they are hardworking or not may systematically differ from how they evaluate whether their friends are hardworking or not. Also, note that the scale for the ‘hardworking’ trait is much smaller than some of the other traits such as gender and local.
4.2 Peer effect measurement
4.2.1 Who are the good students/disruptive peers?
We estimate (3.3) and confirm the presence of a positive peer effect from students who did well in elementary school. We refer to students who did well in elementary school as “good students” and students who did not do well in elementary school as “disruptive peers” therein. Note that since class quantile is a continuous measure, there is a spectrum of “good students” and “disruptive peers”, we use these two terms for easier reference.